News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 15 December 2006 8.00am EDT

The week has been dominated by a series of schedule re-plans due to extensive solar activity.

The loads were restarted at 8:10pm on Dec 9, however were halted by high levels of radiation just prior to entry into the radiation belts at 3:59am on Dec 10. A real-time procedure was completed on Dec 10 to uplink a maneuver-only load to ensure a favorable attitude during earth perigee passage.

Loads were restarted at 9:45pm on Dec 11 however were stopped once more at 5:44pm on Dec 13 due to further solar flares during the day. A real-time procedure was also conducted on Dec 11 to uplink a maneuver-only for earth perigee passage.

During the Dec 13 event, the ACIS instrument experienced a Back End Processor (BEP) reboot caused by the execution of SCS 107 during a bias map computation. The event resulted in a reversion to the launch version of the ACIS flight software. A real-time procedure was completed on Dec 14 to uplink load the ACIS flight software Standard Rev B patches, Optional Patches Rev C and re-start the DEA Housekeeping. This brought the software to the current version. A real-time procedure was executed on Dec 10 to lower the Aspect Camera CCD temperature from -17 to -18 deg C. This is the third of five 1 deg C step reductions planned to mitigate the increasing number of warm pixels.

------------------------------------------
The schedule of targets for the next week will be finalized once the
radiation levels have fallen to acceptable levels.

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports